OpenStudy (anonymous):

help please!! thanks! a satellite, B, is 17,000 miles from the horizon of earth. Earth's radius is about 4,000 miles. Find the approximate distance the satellite is from the Earth's surface.

OpenStudy (anonymous):

@ElonaSushchik

OpenStudy (anonymous):

i will medal

OpenStudy (elonasushchik):

leg 1: 17000 leg 2: 4000 leg 3: x+4000 which allows us to say 17000^2 + 4000^2 = (x+4000)^2 x = sqrt(17000^2 + 4000^2) - 4000 x=?

OpenStudy (elonasushchik):

Can u solve the problem?^

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

but do i do the addition in the first line first? so would i do 17000^2 + 4000 first?

OpenStudy (elonasushchik):

yup..

OpenStudy (anonymous):

i got 289,004,000

OpenStudy (anonymous):

but since there are two 4000 would i subtract each side from 4000?

OpenStudy (anonymous):

wait, i think i got it. would the answer be 13464m?